OpenStudy (anonymous):

Evaluate the expression n= 4 - - 1 7 n- 2 - 4

OpenStudy (jack1):

n - 9/4 = ?? if n = 4/7... is this right?

OpenStudy (jack1):

assuming it is: n - 9/4 = 4/7 - 9/4 = 16/28 - 63 / 28 =...?

OpenStudy (anonymous):

4/7 -9/4 = 16/28-63/28 =- 47/28

OpenStudy (jack1):

@Aliyah13

OpenStudy (anonymous):

@Jack1 that's how I would do it

OpenStudy (jack1):

lol ok id do it completely differently hey ;)